OpenStudy (anonymous):

What is the sum of the geometric sequence 8, –16, 32 … if there are 15 terms?

OpenStudy (anonymous):

a =8, r = -2, n = 15 Sn = a(1-r^n)/(1-r) Sn = 8( 1- (-2)^15) /(1-(-2) Sn = 8/3* (1 - -32768) Sn = 8/3 8 (32769) Sn = 8* 32769/3 =8 *10923 =87,384